# PaperRegister: Boosting Flexible-grained Paper Search via Hierarchical Register Indexing

### v1

@@ Abstract
Paper search is an important activity for researchers, typically involving using a query with description of a topic to find relevant papers. As research deepens, paper search requirements may become more flexible, sometimes involving specific details such as module configuration rather than being limited to coarse-grained topics. However, previous paper search systems are unable to meet these flexible-grained requirements, as these systems mainly collect paper abstracts to construct index of corpus, which lack detailed information to support retrieval by finer-grained queries. In this work, we propose PaperRegister, consisted of offline hierarchical indexing and online adaptive retrieval, transforming traditional abstract-based index into hierarchical index tree for paper search, thereby supporting queries at flexible granularity. Experiments on paper search tasks across a range of granularity demonstrate that PaperRegister achieves the state-of-the-art performance, and particularly excels in fine-grained scenarios, highlighting the good potential as an effective solution for flexible-grained paper search in real-world applications.
Introduction
@@
Paper search is an important and almost everyday activity for researchers (Kuhlthau 1991; Ellis, Cox, and Hall 1993; Hemminger et al. 2007; Case and Given 2016). Typically, this process begins when the user submits a natural-language query describing a topic, then the retrieval system matches this query against the paper corpus and returns the subset of papers with the highest relevance (Wadden et al. 2020; Cohan et al. 2020; Ajith et al. 2024; He et al. 2025). As researchers delve more deeply into their work, the paper search requirements can become increasingly flexible. For example, the view of input query may refer to detailed module configuration or methodological operation, rather than being limited to the level of coarse-grained topic (Mysore et al. 2021; Kang et al. 2024; Wang et al. 2023; Zhang et al. 2025). Therefore, a paper search system supporting queries at flexible levels of granularity is of important value.
Unfortunately, existing paper search systems cannot effectively handle queries across flexible granularity, since they mainly use paper abstracts to construct index of corpus for retrieval (Zheng et al. 2020; Gao et al. 2023; Mackie, Chatterjee, and Dalton 2023; Lei et al. 2024). When a query’s view involves finer‐grained information that does not appear in abstract, they fails to retrieve relevant papers. As shown in Figure 1, the input query “Retrieve papers about jointly training encoder and generator via minimizing the negative marginal log-likelihood without document encoder” specifies a detailed training operation, expecting to find the relevant paper. However, the target cannot be successfully retrieved by traditional paper search systems via abstract-based index, because the abstract does not mention such detailed training operation. Therefore, building a powerful paper search system that supports queries at flexible granularity remains an important open challenge.
@@
Figure 1: PaperRegister can support flexible-grained paper search through the hierarchical register, while traditional method fails due to abstract cannot contain required details.
Figure 2: The overview of PaperRegister, including hierarchical indexing and adaptive retrieval. In offline stage, PaperRegister constructs hierarchical index tree via fine-grained content extracting and bottom-up content aggregating based on hierarchical register schema. In online stage, PaperRegister first identify the views of input query and then conduct view-based matching.
@@
To this end, we propose PaperRegister, which transforms traditional abstract-based index into hierarchical index tree for paper search, thereby supporting queries at flexible granularity. Specially, PaperRegister includes offline hierarchical indexing and online adaptive retrieval. In the offline stage, to construct a hierarchical index tree for corpus, PaperRegister uses a hierarchical register schema, consisted of information nodes at various granularity, in which each node path represents a kind of view. Based on this schema, PaperRegister first extracts fine‐grained contents for each paper, then aggregates contents bottom‐up, layer by layer, yielding hierarchical register of the paper. And then hierarchical index tree is constructed by merging the register of all papers. In the online stage, to adaptively retrieve via the hierarchical index tree, PaperRegister first employs a view recognizer to identify views of input query, which associates to suitable indexes from the hierarchical index tree, then matches query against these indexes to achieve precise paper search.
@@
One key aspect of PaperRegister is the view recognizer, which must be both low latency and high accuracy. To this end, we construct the view recognizer by training a small-scale language model via hierarchical-reward reinforcement learning. Firstly, in order to alleviate the online latency, we adopt a language model with 0.6 billion parameter as base. Then, in order to equip the model with basic identifying capability, we perform supervised fine-tuning (SFT) on it, with query as input and the golden view as label. Furthermore, in order to further enhance accuracy of the view recognizer, considering the hierarchical dependency of views in the hierarchical register schema, we design a hierarchical reward as the reinforcement learning signal, which is calculated by closeness level between the predicted view and golden view in hierarchical register schema. And then we employ this reward on group relative policy optimization (GRPO)  (Shao et al. 2024) training to enhance the view recognizer.
@@
In experiments, we compare PaperRegister with some common-used baselines on paper search tasks across various levels of granularity. The results show PaperRegister achieves state-of-the-art performance, with improvements being more pronounced as query granularity finer, confirming that PaperRegister is a robust solution for challenging flexible-grained paper search tasks. Additionally, detailed ablation and analysis demonstrate the importance of hierarchical register indexing, the effect and performance of view recognizer training, the good compatibility with extra complex paper search frameworks, and the online low-latency advantage compared with several powerful baselines. Code for this work is in https://github.com/Li-Z-Q/PaperRegister.
PaperRegister
@@
As mentioned, existing paper search systems fail to handle flexible-grained queries due to they primarily collect paper abstracts to construct index of corpus, which does not contain detailed information to support queries at finer granularity. To this end, we propose PaperRegister, which supports flexible-grained paper search via hierarchical register indexing. As illustrated in Figure 2, PaperRegister consists of two stages: in the offline stage, a hierarchical index tree for paper corpus is built, and in the online stage, retrieval is performed by adaptively using suitable indexes from the tree.
Task Formulation
@@
The paper search task involved in this work provides a query $q$ as input, with the goal of retrieving $M$ relevant papers $\{p^{(m)}\}_{m=1}^{M}$ based on the index $\mathcal{I}$ of paper corpus $\mathcal{C}$, which can be expressed as the following formula:
@@
For regular paper search, the view of $q$ is typically a coarse-grained topic that user interests, and systems mainly construct index $\mathcal{I}$ based on paper abstracts to match against query $q$. As user’s requirements become more flexible, the view may refer to details such as module configuration and training operation. In this context, since these detailed information typically do not appear in abstract, existing paper search systems cannot perform accurate retrieval. Therefore, we propose PaperRegister to boost flexible-grained paper search, which transforms abstract-based index into hierarchical index tree for paper search, the specific process will be further explained in the following sections.
Offline Hierarchical Indexing
Considering that the granularity of view in queries can be flexible across very fine-grained, moderately fine-grained, to coarse-grained, PaperRegister offline constructs hierarchical index tree for paper corpus to support flexible paper search. Specifically, based on a hierarchical register schema, PaperRegister obtains hierarchical register for each paper through fine-grained content extracting and bottom-up content aggregating. Then the hierarchical index tree is constructed by merging hierarchical register of all papers.
@@
To obtain the hierarchical register for each paper, we design a hierarchical register schema, consisted of information nodes as following:
@@
where $N_{i,j}$ is the $j$-th information node at the $i$-th layer, $n_{i,j}$ is the node name, $c_{i,j}$ is the node content that will be obtained via specific paper text, $N_{i+1,j^{\prime}}$ is a sub-node of $N_{i,j}$, $Z$ is the number of sub-nodes under each node, and $i\in[1,...,L]$, where $L$ is the number of layers in schema.
In this hierarchical register schema, the shallow-layer nodes represent coarse-grained information of the paper, and deep-layer nodes represent finer-grained information. For example, in the i+1i\!+\!1-th layer, $N_{i+1,j\prime}$ and $N_{i+1,j\prime+1}$ respectively denote module configuration and training operation in the paper, which are two kinds of relatively fine-grained information. While their common parent in the $i$-th layer, $N_{i,j}$, denotes method implementation, which is a coarser-grained summary of $N_{i+1,j\prime}$ and $N_{i+1,j\prime+1}$, thereby forming the register schema with hierarchical information nodes.
In addition, since different types of papers have different styles and formats, we design five kinds of hierarchical register schema, corresponding to five types of paper including algorithm innovation, benchmark construction, mechanism exploration, survey, and theory proof. We use a large language model to determine the type of input paper and then assign the corresponding hierarchical register schema. The detailed illustration is presented in the Appendix A.
@@
Based on above hierarchical register schema, in order to obtain the fine-grained contents as much as possible, PaperRegister first perform extracting for each information node at the $L$-th layer based on paper $p^{(m)}$, which can be represented as follows:
@@
where $n_{L,j}$ represents the name of information node $N_{L,j}$ and $\mathcal{M}_{extract}$ is the extracting module for this process.
To achieve accurate extraction, learning from several widely recognized works, in which large language models are proven to possess reliable capabilities for content extraction (Edge et al. 2024; Li et al. 2025c), PaperRegister uses a large language model as the extracting module. Specifically, PaperRegister takes the node name and the text-formatted paper as input, uses instructions to guide in outputting corresponding content for the information node, and leaves it blank if the paper does not include corresponding content. The detailed process and instructions are in Appendix B.
@@
After extracting all the finest-grained contents, in order to obtain the complete hierarchical register, PaperRegister aggregates node contents layer by layer from bottom to top based on the hierarchical register schema, which can be represented as follows:
@@
where layer $i$ is from L1L-1 to 1, thereby obtaining content for each node in the hierarchical register schema for $p^{(m)}$.
To achieve accurate aggregation, as extraction process, PaperRegister uses a large language model as the aggregating module $\mathcal{M}_{aggregate}$. Specifically, PaperRegister takes contents of sub-nodes as input and uses instructions to guide the large language model in summarizing, condensing, and removing details to turn into summary text, thus obtaining the content for upper-layer information node. The detailed process and instructions are presented in the Appendix C.
At this point, PaperRegister obtains node contents at various granular levels to compose the hierarchical register for $p^{(m)}$, and then merges register of all papers in corpus $\mathcal{C}$ to construct hierarchical index tree $\mathcal{I}_{h}$, as shown in following:
$\mathcal{I}_{h}=\{\{{\rm I}_{i,j}\}_{j=1}^{Z}\}_{i=1}^{L}=\{\{\mathcal{M}_{idx}\{c_{i,j}^{(m)}\}_{m=1}^{|\mathcal{C}|}\}_{j=1}^{Z}\}_{i=1}^{L}$ (5)
where $\mathcal{M}_{idx}$ is the indexing module such as BM25 and DPR, each ${\rm I}_{i,j}$ is an index in the hierarchical index tree $\mathcal{I}_{h}$, corresponding to a kind of view for the paper corpus.
Online Adaptive Retrieval
Based on above offline hierarchical indexing, PaperRegister perform online adaptive retrieval via suitable indexes from the hierarchical index tree. Specifically, PaperRegister first identifies views involved in the input query, determining corresponding indexes from the hierarchical index tree, then conduct retrieval by matching query against these indexes.
@@
In order to achieve adaptive retrieval, PaperRegister first uses a view recognizer to identify the view $v_{k}$ in the input query $q$, represented as follows:
@@
where the candidate set of ${v}_{k}$ is all node paths in the hierarchical register schema, $\mathcal{M}_{identify}$ is employed base on a small-scale language model with special training, which will be explained in next section. And to ensure that identifying results can cover the real views of query as more as possible, $\mathcal{M}_{identify}$ uses the beam search strategy (Holtzman et al. 2020) to sample the top-$K$ output views.
@@
After identifying views in the input query, PaperRegister looks up corresponding indexes from the hierarchical index tree $\mathcal{I}_{h}$, as following:
@@
where each ${\rm I}_{k}$ is consisted of $\{c_{k}^{(m)}\}_{m=1}^{|\mathcal{C}|}$ as in Formula 5.
Then these indexes are used to calculate relevance score of input query $q$ and each paper $p^{(m)}$, as following:
@@
where $\mathcal{M}_{rel}$ is relevance module such as BM25 and DPR.
Finally, the retrieved papers with top-$M$ relevance score, $\{p^{(m)}\}_{m=1}^{M}$, are selected as retrieval results for query $q$.
View Recognizer Training
For the PaperRegister system mentioned above, the view recognizer is a key aspect, which must both alleviate the latency and ensure the accuracy of view identifying. To this end, as shown in Figure 3, we use a small-scale language model with 0.6B parameters as the base model, followed by initial supervised fine-tuning (SFT), and then further enhance its view identifying capability by hierarchical-reward group relative policy optimization (GRPO) training.
@@
where $q_{j}$ is a query, $v_{j}$ is the golden view of query, which is a node path in hierarchical register schema. To ensure the stability of training, we make the approximate same number of each view in the training dataset through manual sampling.
@@
To give the small-scale view recognizer basic identifying ability and make it easier for subsequent reinforcement learning training, we first perform SFT by minimizing the following loss:
$\mathcal{L}_{SFT}(\theta)=-\mathbb{E}_{(q,v)\sim\mathcal{D}_{train}}\sum_{t=1}^{|v|}\log\pi_{\theta}(v_{t}|q,v_{<t})$ (10)
where $\pi_{\theta}$ is the model and $v_{t}$ is the $t$-th token of golden view.
@@
To further strengthen the view identifying ability of small-scale view recognizer, we conduct GRPO (Shao et al. 2024) training with hierarchical reward, which is calculated based on the node path’s closeness level of predicted view and golden view in the hierarchical register schema. Specially, considering the hierarchical feature of register schema, different wrong predicted views should not be treated equally. For example, if the golden view’s path in the hierarchical register schema is Abstract-Method-Implementation-Operation, and two predicted views are Abstract-Method-Implementation-Module and Abstract-Experiment-Dataset, although both are wrong, the first one is better than the second one because it is closer to golden view in the hierarchical register schema, suggesting a higher information overlap with the golden view.
Based on this idea, we measure the path overlap between the predicted view $\hat{v}_{j}$ and the golden view $v_{j}$ in the hierarchical register schema, obtaining the hierarchical reward:
@@
where $\mathcal{M}_{overlap}$ returns the overlapping level of two inputs, $|\hat{v}_{j}|$ means the path length of predicted view in the hierarchical register schema and $|v_{j}|$ is that of golden view.
Then this hierarchical reward is used in GRPO training, which is by maximizing the following objective:
$\mathcal{J}_{GRPO}(\theta)={}\mathbb{E}_{\left[q\sim D_{train},\left\{v_{i}\right\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(v\mid q)\right]}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\left|v_{i}\right|}\sum_{t=1}^{\left|v_{i}\right|}\Biggl{\{}$
$\min\Biggl{[}\frac{\pi_{\theta}\left(v_{i,t}\mid q,v_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\left(v_{i,t}\mid q,v_{i,<t}\right)}\hat{A}_{i,t},\text{clip}(\frac{\pi_{\theta}\left(v_{i,t}\mid q,v_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\left(v_{i,t}\mid q,v_{i,<t}\right)},1-\epsilon,1+\epsilon)\hat{A}_{i,t}\Biggr{]}$
$-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\parallel\pi_{\text{ref}}\right]\Biggr{\}},~\text{where}~\hat{A}_{i,t}=\frac{r_{i}-mean(\{r_{1},r_{2},...,r_{G}\})}{std(\{r_{1},r_{2},...,r_{G}\})}$
where $\pi_{\theta_{\text{old}}}$ is initialized by the SFT model, $G$ is the number of each group, $\epsilon$ and $\beta$ are two hyper-parameters.
Method BM25-based Matching DPR-based Matching
@@
Query Paraphrasing
Rewriting 61.9 69.6 58.4 64.2 54.1 59.8 50.9 56.3 76.0 83.2 67.1 72.6 60.4 65.2 55.2 60.9
HyDE 68.9 75.4 60.3 66.3 54.3 60.0 51.5 57.6 76.5 83.4 66.8 72.3 58.9 65.7 56.3 62.4
CSQE 69.0 73.8 59.1 64.4 54.2 58.6 51.3 55.9 77.9 81.7 65.8 71.1 60.2 65.5 55.2 60.9
Paper Splitting
$\text{Chunk}_{avg}$ 58.1 67.6 68.1 74.2 66.6 71.9 65.9 71.6 58.1 67.6 51.1 58.3 46.4 53.4 46.4 52.2
$\text{Chunk}_{max}$ 67.9 75.5 80.0 83.1 81.4 84.7 85.3 87.6 67.9 75.5 79.8 83.1 76.6 80.5 79.0 82.1
$\text{Paragraph}_{avg}$ 29.7 38.3 58.9 66.6 59.8 66.5 58.2 66.2 20.5 22.8 23.3 30.6 23.1 29.8 22.4 29.2
$\text{Paragraph}_{max}$ 64.3 70.9 73.8 78.8 75.4 79.9 80.8 83.9 79.5 85.0 79.2 82.8 76.5 81.4 78.8 81.8
@@
PaperRegister (Ours) 69.7 76.4 89.7 90.9 88.0 89.0 87.5 88.7 81.0 87.1 84.1 87.1 79.9 82.5 80.8 82.9
Table 1: Main experimental results on paper-searching tasks across various granularity, where the granularity is coarse-to-fine from LitSearch to F.g.Search-3. The table shows that PaperRegister is a powerful solution to addressing flexible-grained paper search and the advantages of PaperRegister become more pronounced at finer granular queries.
Experiments
Experimental Settings

### v2

@@ Abstract
As researchers delve more deeply into their work, paper search requirements may become more flexible, sometimes involving specific details such as module configuration rather than being limited to coarse-grained topics. However, previous paper search systems are unable to meet these flexible-grained requirements, as previous systems mainly collect paper abstract to construct corpus index, which lacks detailed information to support retrieval by some finer-grained queries. In this work, we propose PaperRegister, which transforms traditional abstract-based index into a hierarchical index tree, thereby supporting queries at flexible granularity. Experiments on paper search tasks across a range of granularity demonstrate that PaperRegister achieves the SOTA performance, and particularly excels in the fine-grained scenarios, highlighting good potential as an effective solution for flexible-grained paper search in real-world applications. https://github.com/Li-Z-Q/PaperRegister.
PaperRegister: Boosting Flexible-grained Paper Search via
Hierarchical Register Indexing
Zhuoqun Li1,2, Xuanang Chen1, Hongyu Lin1, Yaojie Lu1, Xianpei Han1, Shanshan Jiang3, Bin Dong3, Le Sun1 1Chinese Information Processing Laboratory, Institute of Software, Chinese Academy of Sciences 2University of Chinese Academy of Sciences 3Ricoh Software Research Center Beijing, Ricoh Company, Ltd. {lizhuoqun2021,chenxuanang,hongyu,luyaojie,xianpei,sunle}@iscas.ac.cn
1 Introduction
@@
Paper search is an important and almost everyday activity for researchers kuhlthau1991inside; ellis1993comparison; hemminger2007information; case2016looking. Typically, this process begins when user submits a natural-language query describing a topic, and then retrieval system matches this query against paper corpus and returns subset of papers with the highest relevance wadden2020fact; cohan2020specter; ajith2024litsearch; he2025pasa. As researchers delve more deeply into their work, paper search requirements can become increasingly flexible. For example, the view of query may refer to detailed module configuration or methodological operation, rather than being limited to the level of coarse-grained topic mysore2021csfcube; kang2024taxonomy; wang2023scientific; zhang2025scientific. Therefore, a paper search system supporting queries at flexible granularity is of important value.
@@
Figure 1: PaperRegister supports flexible-grained paper search via hierarchical register, while traditional method fails due to abstract cannot contain required details.
Unfortunately, existing paper search systems cannot effectively handle queries across flexible granularity, since they mainly use paper abstract to construct corpus index for retrieval zheng2020bert; gao2023precise; mackie2023generative; lei2024corpus. When the view of query involves finer‐grained information that does not appear in abstract, they fail to retrieve relevant papers. As shown in Figure 1, the query “Retrieve papers about jointly training encoder and generator via minimizing negative marginal log-likelihood without doc encoder” focuses on a detailed training operation. For this query, the target papers cannot be successfully retrieved by traditional abstract-based index, because paper abstract does not mention such a detailed training operation. Therefore, building a paper search system supporting queries at flexible granularity remains a valuable and unresolved challenge.
@@
To this end, we propose PaperRegister, which transforms traditional abstract-based index into a hierarchical index tree, thereby supporting queries at flexible granularity. Specially, PaperRegister includes offline hierarchical indexing and online adaptive retrieval. In the offline stage, to construct the hierarchical index tree for paper corpus, PaperRegister uses a hierarchical register schema, which is consisted of information nodes at various granularity, and each node path represents a kind of view. Based on this schema, PaperRegister first extracts fine‐grained node contents, then aggregates node contents bottom‐up, layer by layer, yielding hierarchical register for each paper. And then the hierarchical index tree is constructed by merging register of all papers. In the online stage, to adaptively retrieve via the hierarchical index tree, PaperRegister first employs a view recognizer to identify the views of query, which associates to suitable indexes from the tree, and then matches query against these indexes to achieve the precise paper search.
@@
One key aspect of PaperRegister is the view recognizer, which must be both low-latency and high-accuracy. To this end, we construct the view recognizer by training a small-scale language model via hierarchical-reward reinforcement learning. Firstly, to alleviate online latency, we adopt a language model with 0.6 billion parameters as base. Then, to equip the model with basic identifying ability, we perform supervised fine-tuning, where query is the input and golden view is the label. Furthermore, to enhance accuracy of view recognizer, considering the hierarchical dependency of nodes in register schema, we design a hierarchical reward as reinforcement learning signal, which is calculated by the closeness level between predicted view and golden view in hierarchical register schema. And then we employ this reward on group relative policy optimization (GRPO)  Shao2024DeepSeekMathPT to enhance the capability of view recognizer.
Figure 2: PaperRegister includes hierarchical indexing and adaptive retrieval. Offline, PaperRegister constructs hierarchical index tree via fine-grained content extracting and bottom-up content aggregating based on a hierarchical register schema. Online, PaperRegister first identify views of query and then conduct view-based matching.
@@
In experiments, we compare PaperRegister with several common-used methods on paper search, across various levels of granularity. Results show PaperRegister achieves the SOTA performance, with the improvement being more pronounced as query granularity is finer, confirming that PaperRegister is a robust solution for the challenging flexible-grained paper search task. The main contributions of this work can be summarized as follows:
  • We construct PaperRegister, which can support paper search queries at flexible granularity via the hierarchical register indexing.
  • We design hierarchical-reward reinforcement learning, which can train a powerful and low-latency view recognizer for PaperRegister.
  • We conduct extensive experiments and analysis, which show PaperRegister is an advanced system for flexible-grained paper search.
2 PaperRegister
@@
Existing paper search systems fail to handle queries at flexible granularity due to they primarily collect paper abstract to construct index of corpus, which does not contain detailed information to support queries at finer granularity. To this end, we propose PaperRegister, which can support flexible-grained paper search via hierarchical register indexing. As illustrated in Figure 2, in the offline stage, a hierarchical index tree is built for paper corpus, and in the online stage, adaptive retrieval is performed by selecting and using suitable indexes from the tree.
2.1 Task Formulation
@@
The task involved in this work provides a query $q$ as input, with the goal of retrieving $M$ relevant papers $\{p^{(m)}\}_{m=1}^{M}$ based on the index $\mathcal{I}$ of paper corpus $\mathcal{C}$, which can be expressed as follows:
@@
For regular paper search, view of query $q$ is typically a coarse-grained topic that user is interested in, and systems mainly construct index $\mathcal{I}$ based on paper abstract to match against $q$. As user’s requirements become more flexible, the view may refer to details such as module configuration and training operation. In this context, since this detailed information typically do not appear in abstract, existing paper search systems cannot retrieve accurately.
2.2 Offline Hierarchical Indexing
Considering the granularity of query view can be flexible across very fine-grained, moderately fine-grained, to coarse-grained, PaperRegister offline constructs a hierarchical index tree for paper corpus to support flexible paper search. Specifically, based on a hierarchical register schema, PaperRegister obtains hierarchical register for each paper through fine-grained content extracting and bottom-up content aggregating, and then hierarchical index tree is constructed by merging register of all papers.
@@ 2.2 Offline Hierarchical Indexing
To obtain the hierarchical register for each paper, we first design a hierarchical register schema, which consists of information nodes as the following formula:
@@
where $N_{i,j}$ is the $j$-th information node at the $i$-th layer, $n_{i,j}$ is node name, $c_{i,j}$ is expected node content from specific paper, $N_{i+1,j^{\prime}}$ is a sub-node of $N_{i,j}$, $Z$ is number of sub-nodes under each node, and $i\in[1,...,L]$, $L$ is number of layers in schema.
In hierarchical register schema, shallow-layer nodes represent coarse-grained information, and deep-layer nodes represent finer-grained information. For example, in the i+1i\!+\!1-th layer, $N_{i+1,j\prime}$ and $N_{i+1,j\prime+1}$ respectively denote module configuration and training operation of paper, which are two kinds of relatively fine-grained information. While their common parent in the $i$-th layer, $N_{i,j}$, denotes method implementation, which is a coarser-grained summary of $N_{i+1,j\prime}$ and $N_{i+1,j\prime+1}$, thereby forming schema with hierarchical-dependency nodes.
In addition, since different types of papers have different styles and formats, we design five kinds of schema, corresponding to five paper types, and use a large language model to determine the type of each paper. Details are presented in Appendix A.
@@
Based on the above hierarchical register schema, in order to obtain fine-grained contents as much as possible, PaperRegister first extracts content for each information node at the $L$-th layer from paper $p^{(m)}$, which can be represented as the following formula:
@@
where $n_{L,j}$ represents name of information node $N_{L,j}$ and $\mathcal{M}_{extract}$ is the extracting module.
To achieve accurate extraction, learning from several widely recognized works, in which large language models are proven to possess reliable capabilities for content extraction tasks edge2024local; li2025deepsolution, PaperRegister employs a large language model as extracting module. Specifically, PaperRegister takes the node name and text-formatted paper as input, uses instructions to guide in outputting corresponding content for the information node, and leaves it blank if the paper does not include corresponding content. The detailed process and instructions are in the Appendix B.
@@
After extracting all the finest-grained contents, in order to smoothly obtain a complete hierarchical register, PaperRegister aggregates node contents layer by layer from bottom to top based on the hierarchical register schema, represented as following formula:
@@
where the layer $i$ is from L1L-1 to 1, thereby obtaining content for each information node in the hierarchical register schema for the paper $p^{(m)}$.
To achieve accurate aggregation, like extraction process, PaperRegister uses a large language model as the aggregating module $\mathcal{M}_{aggregate}$. Specifically, PaperRegister takes contents of sub-nodes as input and uses instructions to guide large language model in summarizing, condensing, and removing details to turn into summary text, thus obtaining content for upper-layer information node. The detailed process and instructions are in Appendix C.
At this point, PaperRegister obtains node contents at various granular levels to compose the hierarchical register for $p^{(m)}$. And then merges register of all papers in corpus $\mathcal{C}$ to construct hierarchical index tree $\mathcal{I}_{h}$, as shown in following formula:
$\mathcal{I}_{h}=\{\{{\rm I}_{i,j}\}_{j=1}^{Z}\}_{i=1}^{L}=\{\{\mathcal{M}_{idx}\{c_{i,j}^{(m)}\}_{m=1}^{|\mathcal{C}|}\}_{j=1}^{Z}\}_{i=1}^{L}$
where $\mathcal{M}_{idx}$ is the indexing module such as BM25 and DPR, each ${\rm I}_{i,j}$ is an index in hierarchical index tree, corresponding to a kind of view for corpus.
2.3 Online Adaptive Retrieval
Based on the offline hierarchical indexing above, PaperRegister perform online adaptive retrieval via suitable indexes from hierarchical index tree. Specifically, PaperRegister first identifies views involved in the query, determining corresponding indexes from hierarchical index tree, then conduct retrieval by matching query against these indexes.
@@ 2.3 Online Adaptive Retrieval
To achieve adaptive retrieval, PaperRegister first uses a view recognizer to identify the view $v_{k}$ in query $q$, represented as follows:
@@
where the candidate set of ${v}_{k}$ is all node paths in the hierarchical register schema, $\mathcal{M}_{identify}$ is employed base on a small-scale language model with special training, which will be explained in next section. And to ensure that identifying results can cover the real views of query as more as possible, $\mathcal{M}_{identify}$ uses the beam search strategy Holtzman2020The to sample the top-$K$ output views.
@@
After identifying views in query, PaperRegister looks up corresponding indexes from hierarchical index tree $\mathcal{I}_{h}$, as follows:
@@
where ${\rm I}_{k}$ is consisted of $\{c_{k}^{(m)}\}_{m=1}^{|\mathcal{C}|}$ as Formula 5.
Then these indexes are used to calculate relevance score of input query $q$ and each paper $p^{(m)}$  :
@@
where $\mathcal{M}_{rel}$ is relevance module such as BM25.
Finally, papers with top-$M$ relevance score, $\{p^{(m)}\}_{m=1}^{M}$, are selected as final results for query.
3 View Recognizer Training
For PaperRegister system, view recognizer is a key module, which must both alleviate latency and ensure accuracy. To this end, as in Figure 3, we use a small-scale language model with 0.6B parameters as base, first do supervised fine-tuning (SFT), and then further enhance capability by hierarchical-reward group relative policy optimization (GRPO).
@@ 3 View Recognizer Training
where $q_{j}$ is a query, $v_{j}$ is golden view of query, which is node path in hierarchical register schema.
@@
To give small-scale view recognizer basic identifying ability and make it easier for subsequent reinforcement learning, we first perform SFT by minimizing following loss:
$\mathcal{L}_{SFT}(\theta)=-\mathbb{E}_{(q,v)\sim\mathcal{D}_{train}}\sum_{t=1}^{|v|}\log\pi_{\theta}(v_{t}|q,v_{<t})$
where $\pi_{\theta}$ is model, $v_{t}$ is the $t$-th golden view token.
@@
To strengthen the capability of small-scale view recognizer, we conduct GRPO Shao2024DeepSeekMathPT with hierarchical reward, which is calculated via the closeness level of predicted view and golden view in hierarchical register schema. Specially, considering the hierarchical-dependency feature of register schema, different wrong predicted views should not be treated equally. For example, the golden view of query is represented as Abstract-Method-Implementation-Operation in schema, and two predicted views are Abstract-Method-Implementation-Module and Abstract-Experiment-Dataset respectively. Although both are wrong, the first one is better than the second because it is closer to golden view in hierarchical register schema, suggesting a higher information overlap with golden view.
Based on this, we measure path overlap between predicted view $\hat{v}_{j}$ and golden view $v_{j}$ in hierarchical register schema, obtaining hierarchical reward:
@@
where $\mathcal{M}_{overlap}$ returns overlap level of inputs, $|\hat{v}_{j}|$ means path length of predicted view in hierarchical register schema, and $|v_{j}|$ is that of golden view.
Then this hierarchical reward is used in GRPO, which is by maximizing the following objective:
$\begin{aligned} \mathcal{J}_{GRPO}(\theta)={}&\mathbb{E}_{\left[q\sim D_{train},\left\{v_{i}\right\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(v\mid q)\right]}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\left|v_{i}\right|}\sum_{t=1}^{\left|v_{i}\right|}\Biggl\{\\ &\min\Biggl[\frac{\pi_{\theta}\left(v_{i,t}\mid q,v_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\left(v_{i,t}\mid q,v_{i,<t}\right)}\hat{A}_{i,t},\text{clip}(\frac{\pi_{\theta}\left(v_{i,t}\mid q,v_{i,<t}\right)}{\pi_{\theta_{\text{old}}}\left(v_{i,t}\mid q,v_{i,<t}\right)},1-\epsilon,1+\epsilon)\hat{A}_{i,t}\Biggr]\\ &-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\parallel\pi_{\text{ref}}\right]\Biggr\},~\text{where}~\hat{A}_{i,t}=\frac{r_{i}-mean(\{r_{1},r_{2},...,r_{G}\})}{std(\{r_{1},r_{2},...,r_{G}\})}\end{aligned}$
where $\pi_{\theta_{\text{old}}}$ is initialized by SFT model, $G$ is number of each group, $\epsilon$ and $\beta$ are hyper-parameters.
Method BM25-based Paper Search DPR-based Paper Search
@@
Query Enhancing
Rewriting ma2023query 61.9 69.6 58.4 64.2 54.1 59.8 50.9 56.3 76.0 83.2 67.1 72.6 60.4 65.2 55.2 60.9
HyDE gao2023precise 68.9 75.4 60.3 66.3 54.3 60.0 51.5 57.6 76.5 83.4 66.8 72.3 58.9 65.7 56.3 62.4
CSQE lei2024corpus 69.0 73.8 59.1 64.4 54.2 58.6 51.3 55.9 77.9 81.7 65.8 71.1 60.2 65.5 55.2 60.9
Multi-field Indexing
$\text{Chunk}_{avg}$ li2025multifield 58.1 67.6 68.1 74.2 66.6 71.9 65.9 71.6 49.7 58.6 51.1 58.3 46.4 53.4 46.4 52.2
$\text{Chunk}_{max}$ li2025multifield 67.9 75.5 80.0 83.1 81.4 84.7 85.3 87.6 72.6 79.7 79.8 83.1 76.6 80.5 79.0 82.1
$\text{Paragraph}_{avg}$ li2025multifield 29.7 38.3 58.9 66.6 59.8 66.5 58.2 66.2 20.5 22.8 23.3 30.6 23.1 29.8 22.4 29.2
$\text{Paragraph}_{max}$ li2025multifield 64.3 70.9 73.8 78.8 75.4 79.9 80.8 83.9 79.5 85.0 79.2 82.8 76.5 81.4 78.8 81.8
@@
PaperRegister (Ours) 69.7 76.4 89.7* 90.9* 88.0* 89.0* 87.5* 88.7* 81.0* 87.1* 84.1* 87.1* 79.9* 82.5* 80.8* 82.9*
Table 1: Main results on paper search across various granularity, where granularity is coarse-to-fine from LitSearch to F.g.Search-3. Results shows that PaperRegister is an effective system for flexible-grained paper search and the advantage of PaperRegister become more pronounced at finer granular tasks. The * indicates statistical significance (p < 0.05) compared with the best baseline. And experiment results by more metrics are reported in Appendix E.
4 Experiments
4.1 Experimental Settings